TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Created by Main.AntonioVergara on 2007-06-26 - 20:06

Powering Tests Sector 7-8 - Tuesday 26.06.2007

PDF Progress_PoweringS78_070627.pdf:

Test Report from Tuesday 26.06.2007

  • RB.A78
    • PIC2 - Test completed OK
    • During the whole day, the noise read by the bus-bar quench detector of the circuit tripped several times. The source of the noise (period of about 20 minutes) has not been identified yet.
    • PLI1 (760A)
      • PLI1.1 - Current cycle - Test OK
      • PLI1.2 - Energy extraction discharge (quench to a current lead) - Test OK
      • PLI1.3 - Powering Failure from PC - Test OK

  • RQ5.L8: Commissioned up to PLI3 using the sequencer - Tests OK

  • 60A closed orbit correctors - P2N completed for one corrector using new automated procedure

  • RQF.A78 - Type test completed - Test OK

  • 600A correctors
    • RCO.A78 and RQT12.L8B1 - PCS completed - Test OK

Test Plan for Wednesday 27.06.2007

  • Show goes from 7:00 to 22:00

  • RB.A78
    • PLI1.4 - FPA from PIC
    • PLI1.5 - Quench heater discharge on A28.R7
    • PLI2 (2000A)
      • PLI2.1 - Current cycle
      • PLI2.2 - Energy extraction discharge (quench to a current lead)
      • PLI2.2b - SPA from PIC
      • PLI2.3 - Quench heater discharge on LBBRA.13L8 (magnet 3002) and LBARA.13R7 (magnet 1054)

  • RQF.A78 and RQD.A78 (once RB.A78 tests are completed)
      • RQF and RQD ramped up together to 6500A. Quench on 15L8
      • RQF and RQD ramped up together to 6500A. Quadruple quench 20L8, 19L8, 19R7 and 22R7 .

  • Simmultaneous powering of RQF.A78, RQD.A78 and RB.A78.

  • 600A Correctors - More circuits will be released in the afternoon by QPS for PCS tests

  • 60A Closed Orbit Correctors
    • P2N will continue
    • Type tests for some converters in the evening after quenches in the main circuits

  • Matching Section
    • Reproduction of the quench event of May 31st
    • Last commissioning step of RQ5.L8 via the sequencer

  • AB/PO (David) will prepare a list of the activities to be carried out in the conveters already commissioned before the warm-up

Edit, attach file or add comment

Created by Main.MirkoPojer on 2007-06-26 - 19:00

Beam Dump System - Installation and Commissioning - 19/06/2007

PDF RAT_20070619.pdf:


Edit, attach file or add comment

Created by Main.UnknownUser on 2007-06-26 - 18:47

Diagnosis of MB.A45 circuit shorted to ground - Courtesy of D.Bozzini & J.P.Tock

This morning MCS has open the diode box of the dipole 3186. AT/MEL has connected an ohmmeter during the operation to follow the status of the short. At the time of opening the big flange of the diode box the resistance to ground started to change (increase). After the opening it could be possible to first see a quite big quantity of dust on the diode metallic box and second some big pieces of welding residuals.

After having disconnected the diode at the level of the half moons the short disappeared (infinite resistance) and in at the same time some debris felled down from the top part of the half moons. As you can see in the attached pictures the size of one of this residual fits perfectly between the live part and the metallic body of the bus bar duct. In addition in one picture you will see a carbon trace on the G10 insulating piece. This could be provoked by the current of 1A we injected to better localize the fault position flowing through the debris supported onto the G10 piece.

The system will be set back to the requested pressure condition (6 bars in the magnets, 2.2 bars in the DFBA's - which should be achieved by Tuesday 3rd July), when the ELQA-TP4B qualification of this faulty circuit will be repeated.

JPEG IMG_0871.jpg:

JPEG IMG_0933.jpg:

JPEG IMG_0909.jpg:

JPEG IMG_0882.jpg:


Edit, attach file or add comment

Created by Main.BorisBellesia on 2007-06-26 - 17:48

Short circuit test UA23 - 26th June 2007

PDF RAT_20070626.pdf:


Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login